OpenStudy (anonymous):

Graph the expression sec2 x - sec2 x sin2 x on your calculator. Determine what constant or single function is equivalent to the given expression.

OpenStudy (cwrw238):

are the 2\s powers?

OpenStudy (anonymous):

I got it. Nevermind.

OpenStudy (cwrw238):

if so = sec^2 x( 1 - sin^x) 1- sin^2 x = cos^2 x so we have sec^2x * cos^2 x cos^2x = ------- cos^2 x

OpenStudy (cwrw238):

= 1

OpenStudy (cwrw238):

agreed?

OpenStudy (anonymous):

Yep

OpenStudy (anonymous):

\[\sec^2x-\sec^2x~\sin^2x\]\[\left({1\over\cos^2x}\right)-\left({1\over\cos^2x}\right)~\sin^2x\]\[\left({1\over\cos^2x}\right)-\left({\sin^2x\over\cos^2x}\right)\]\[~~{1-\sin^2x\over\cos^2x}\]\[~~~{\cos^2x\over\cos^2x}\]\[~~~=1\]